Calculate the in the [H+] following solution.

[OH-}=6.7x10^-4 M at 25 °C = M

Kw= 1.0x10^-14 at 25 °C = M

[H+]=_______ M
Answer #1

Kw = [H+][OH-]

[H+] = Kw / [OH-]

[H+] = 1.0×10^-14 / 6.7×10^-4

[H+] = 1.49×10^-11

[H+] = 1.5×10^-11 M